OpenStudy (anonymous):

Curvature of the helix r(t)=a cost i +a sint j +bt k (a,b>= 0) and curvature K= a/(a^2 +b^2) . What is the largest value of K for a given value of b

OpenStudy (amistre64):

\[k=\frac{|r'xr''|}{|r'|}\] or am i missing a ^2 underneath?

OpenStudy (amistre64):

^3 in the denom

OpenStudy (amistre64):

r(t)=<a cost, sint,bt > r'(t)=<-a sint, cost,b > r''(t)=<-a cost, -sint, 0>

OpenStudy (anonymous):

for curvature we are using the formula K= 1/IVI * ldT/dtl

OpenStudy (amistre64):

r' <-a sint, cost,b > xr''<-a cost, -sint, 0> -------------------- x = -b sint -y = (-ab cost) z = a (sin^2t + cos^2t) r'xr'' = <-b sint, ab cost, a>

OpenStudy (amistre64):

yeah, I believe that simplifies to this |r'xr''|/|r'|^3

OpenStudy (anonymous):

ok

OpenStudy (dominusscholae):

Don't you just have to maximize the curvature they give already by deriving with respect to b? Because from what I can tell, finding the curvature is going to give you just what is already given.

OpenStudy (amistre64):

|r'xr''| = sqrt( b^2 sin^2t + (ab)^2 cos^2t + a^2)

OpenStudy (anonymous):

yes i am trying to find the maximum value of K with respect to b @ dominusscholae

OpenStudy (amistre64):

i gotta get to class, so yall have fun with it ;)

OpenStudy (anonymous):

thank amistre

OpenStudy (anonymous):

thanks*

OpenStudy (anonymous):

dominusscholae can you help me with this question

OpenStudy (dominusscholae):

Yeah. From what I can tell, you just derive K with respect to b to find the critical point. We are just assuming a is constant so that we find the correct b value at which K is maximum.